OpenStudy (anonymous):

If 1.2 kilograms of rust form on a bridge in five days, what should be the rate of reaction in grams per hour? (Recall that a bar over the number means that zero is significant.)

thomaster (thomaster):

1.2 KG = 1200 gram 1200 gram per 5 days 5 days = 5*24=120 hours so 1200 grams or rust in 120 hours how many gram is that per hour?

OpenStudy (anonymous):

10g/hr?

OpenStudy (anonymous):

6 g/hr

thomaster (thomaster):

eee i made a mistake :P \(\large\dfrac{1200}{120}=10~g/hr\) you were right :)